OpenStudy (anonymous):

Use elimination to solve the system. 3x-12=5y 51-5y=4x

OpenStudy (anonymous):

Why not use substitution? it is easier with subs.

OpenStudy (anonymous):

3x-12=y 3x=y+12 x=(y+12)/3 51-5y=4x 51-5y = 4(y+12)/3 3(51-5y)=4(y+12) 153-15y = 4y+48 -15y-4y = 48-153 -19y = -105 y = -105/-19 = 5

OpenStudy (anonymous):

the first eq states what 5t equals, then the second one uses 5y

OpenStudy (anonymous):

*y, not t

OpenStudy (anonymous):

Kush, you said: 3x-12=y 3x=y+12 x=(y+12)/3 but it isn't 3x-12=y it is 5y

OpenStudy (anonymous):

3x-12=5y 51-5y=4x 51-(3x-12)=4x 51-3x+12=4x 63 = 7x x = 9 y = 3

OpenStudy (anonymous):

I like subs too but i have to show work using elimination, which i get stumped on a lot

OpenStudy (anonymous):

\[3x-12=5y\] \[3x=5y+12\] \[x=\frac{5y+12}3\] \[51-5y=4x\] \[51-5y=\frac{4(5y+12)}3\] \[3(51-5y)=4(5y+12)\] \[153-15y=20y+48\] \[-15y-20y=48-153\] \[-35y = -105\] \[y=\frac{-105}{-35}\] \[y=3\]\[x=9\]

OpenStudy (anonymous):

3x-12=5y 51-5y=4x ok, first set them up so like terms are in line vertically: 3x - 5y = 12 4x +5y = 51

OpenStudy (anonymous):

then, since there is a +5y and a -5y, add the equations' terms: 3x - 5y = 12 4x +5y = 51 7x + 0 = 63 You can take it from there

OpenStudy (anonymous):

i think there should be so much explaination needed .....tonks shannon where r u stuck ? tell us we would clear ur doubt !